OpenStudy (anonymous):

Multiply x^(1/2) times x^(3/7)

OpenStudy (anonymous):

answer choices:

OpenStudy (anonymous):

I picked D?

OpenStudy (souvik):

\[\huge{x ^{a}\times x ^{b}=x ^{(a+b)}}\]

OpenStudy (anonymous):

but what if theyre fractions?

OpenStudy (souvik):

\[x^(\frac{ 1 }{ 2}+\frac{ 3 }{ 7 })\]

OpenStudy (anonymous):

oh so its A?

OpenStudy (anonymous):

thanks!

OpenStudy (souvik):

np:)